Accounts receivable changes with bad debts - A firm is evaluating an accounts receivable change that would increase bad debts from2% to 3% of sales. Sales are currently 40,000 units, the selling price is $20 per unit, and the variable cost per unit is $15. As a result of the proposed change sales are forecast to increase to 70,000 units.

a. what are the bad debts in dollars currently and under proposed change?

b. calculate the cost of the marginal bad debts to the firm?

c. ignoring the additional profit contribution from increased sales, if the proposed change saves $3,500 and causes no change in the average investment in accounts receivable, would you recommend it?

d. considering all changes and benefits, would you recommend the proposed change?

e. compare and discuss your answers in parts c and d?
Answer #1

Solution - Given that current Sales - 40,000 unity The Selling price = $20 per unit Variable cost = $15 per unit changing sald) Increas in sale units = 70,000-40,000 = 30,000 units contribution spavc=20-15 = 5P.U Contribution value = 30,000 x 5 = 450